OpenStudy (anonymous):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 15(1.04)n Part A: When the scientist concluded his study, the height of the plant was approximately 20.53 cm. What is a reasonable domain to plot the growth function? Part B: What does the y-intercept of the graph of the function f(n) represent? Part C: What is the average rate of change of the function f(n) from n = 1 to n = 5, and what does it represent?

OpenStudy (unklerhaukus):

which variable do you think the domain will be ?

OpenStudy (anonymous):

20

OpenStudy (unklerhaukus):

twenty is not a variable

OpenStudy (anonymous):

n

OpenStudy (unklerhaukus):

yes 'the number of days' : n , is the independent variable, that goes on the x axis

OpenStudy (unklerhaukus):

what domain for n is considered in the experiment ?

OpenStudy (anonymous):

I dont know please help me

OpenStudy (unklerhaukus):

I guess the scientist began on day zero ,

OpenStudy (anonymous):

Yes i think so That the starting point

OpenStudy (unklerhaukus):

He said that he finished when the plant was 20.53 [cm], For what values of n does the height function equal 20.53? f(n) =15(1.04)^n f(End of Domain)= 20.53

OpenStudy (unklerhaukus):

If we call the end point E, f(E) = 15(1.04)^E = 20.53 Can you solve this for E ?